Exercise 9-10

During 2020, Bramble Furniture Company purchases a carload of wicker chairs. The manufacturer sells the chairs to Bramble for a lump sum of $35,910 because it is discontinuing manufacturing operations and wishes to dispose of its entire stock. Three types of chairs are included in the carload. The three types and the estimated selling price for each are listed below.

Type

No. of Chairs

Estimated Selling
Price Each

Lounge chairs

240 $90

Armchairs

180 80

Straight chairs

420 50

During 2020, Bramble sells 120 lounge chairs, 60 armchairs, and 72 straight chairs.

What is the amount of gross profit realized during 2020? What is the amount of inventory of unsold straight chairs on December 31, 2020? (Round cost per chair to 2 decimal places, e.g. 78.25 and final answer to 0 decimal places, e.g. 5,845.)

Gross profit realized during 2020

$enter a dollar amount

Amount of inventory of unsold straight chairs

$enter a dollar amount
Answer #1

A

B

C = AXB

D = C/total of C

E

F = D X E

G = F/A

Particulars

No. of
chairs

Sales price
per chair

Total sales
price

Relative to
sales price

Total cost

cost allocated
to chair

cost per chair

Lounge chairs

240

90

21600

0.37894737

35910

13608

56.7

Arm chairs

180

80

14400

0.25263158

35910

9072

50.4

Straight chairs

420

50

21000

0.36842105

35910

13230

31.5

Totals

57000

Particulars

No. of
chairs

cost per chair

Cost of sold
chairs

Selling price

Sales value

Gross profit =
Sales -cost

Lounge chairs

120

56.7

6804

90

10800

3996

Arm chairs

60

50.4

3024

80

4800

1776

Straight chairs

72

31.5

2268

50

3600

1332

Totals

12096

19200

7104

Gross profit = $ 7104

Amount of inventory of unsold straight chairs = (420-72) X31.5 =$ 10962